# Zhaoyong Huang

### 05 Mar 2026 v1

@@ 1 Introduction
In Section 4, we apply the results obtained in Section 3 to module categories. The study of generalized tilting modules, which were usually called Wakamatsu tilting modules, was initiated by Wakamatsu [40]. It follows from [42, Corollary 3.2] that $_{R}C$} is Wakamatsu tilting if and only if $C_{S}$} is Wakamatsu tilting with $R=\mathop{\rm End}\nolimits(C_{S})$, if and only if $_{R}C_{S}$ is semidualizing bimodule in the sense of [16]. Many authors have studied the properties of Wakamatsu tilting modules and semidualizing bimodules and related modules, see [1, 3], [12], [15, 16], [19][22], [26], [28], [35][42] and references therein.
Let $R$ be an arbitrary ring and $_{R}C$ a Wakamatsu tilting module with $R=\mathop{\rm End}\nolimits(C_{S})$. We use $\mathcal{P}_{C}(R)$ and $\mathcal{F}_{C}(R)$ to denote the classes consisting of $C$-projective and $C$-flat left $R$-modules, respectively, and use $\mathcal{I}_{C}(S)$ and $\mathcal{FI}_{C}(S)$ to denote the classes consisting of $C$-injective and $C$-FP-injective left $S$-modules, respectively (see Section 2.2 for the definitions of these modules). As an application of Theorem 1.1, we get the following result.
@@
      The category of left $R$-modules is right $n$-$\mathcal{P}_{C}(R)$-Gorenstein.
@@
      The category of left $S$-modules is left $n$-$\mathcal{I}_{C}(S)$-Gorenstein.
@@
The Auslander class $\mathcal{A}_{C}(S)$ and Bass class $\mathcal{B}_{C}(R)$ with respect to a semidualizing bimodule $_{R}C_{S}$} (equivalently, a Wakamatsu tilting module $_{R}C$ with $R=\mathop{\rm End}\nolimits(C_{S})$) originated from [12] are now collectively known as Foxby classes, and they are linked together by Foxby equivalence [12, 16] and duality pairs [20]. These two classes of modules play a crucial role in studying the homological theory related to semidualizing bimodules and Wakamatsu tilting modules, see [1, 12, 15, 16], [19][22], [26, 28], [35][42] and references therein.
@@
    The category of left $R$-modules is left $n$-$\mathcal{P}_{C}(R)$-Gorenstein.
@@
    The category of left $S$-modules is right $n$-$\mathcal{I}_{C}(S)$-Gorenstein.
@@ 2.1 Relative homological dimensions and one-sided Gorenstein subcategories
$\text{in}\ \mathscr{A}\ \text{with all}\ X_{i}\in\mathscr{X}\},$
@@
$\text{in}\ \mathscr{A}\ \text{with all}\ X^{i}\in\mathscr{X}\},$
@@ 2.2 Wakamatsu tilting modules
Generalized tilting modules are usually called Wakamatsu tilting modules, see [3, 28]. Let $C$ be a left $R$-module and $S=\mathop{\rm End}\nolimits(_{R}C)$. It follows from [42, Corollary 3.2] that $_{R}C$} is Wakamatsu tilting if and only if $C_{S}$} is Wakamatsu tilting with $R=\mathop{\rm End}\nolimits(C_{S})$, if and only if $_{R}C_{S}$ is semidualizing bimodule in the sense of [1, 16]. Typical examples of Wakamatsu tilting modules include the free module of rank one and the dualizing module over a Cohen-Macaulay local ring. For more examples of Wakamatsu tilting modules, the reader is referred to [16, 37, 41].
@@ 2. $(b)$
    in $\mathscr{A}$ with all $U_{i}\in\mathscr{U}$.
@@
in $\mathscr{A}$ with all $U_{i}\in\mathscr{U}$ is $\mathop{\rm Hom}\nolimits_{\mathscr{A}}(\mathscr{U},-)$-exact, and the assertion follows.
@@
    in $\mathscr{A}$ with all $E^{i}\in\mathscr{E}$.
@@ 3.2 One-sided Gorenstein categories
in $\mathscr{A}$ with all $G_{i}\in r\mathcal{G}(\mathscr{C},\mathscr{D})$. For any $D\in\mathscr{D}$, applying the functor $\mathop{\rm Hom}\nolimits_{\mathscr{A}}(-,D)$ to the above exact sequence yields
@@
It is easy to see that $(4)\Longrightarrow(1)$ and $(5)\Longrightarrow(3)$.
@@
    $\mathscr{A}$ is right $n$-$\mathscr{C}$-Gorenstein ((that is, $\mathop{\rm pd}\nolimits\mathcal{I}(\mathscr{A})\leq n$ and $\mathop{\rm id}\nolimits\mathscr{D}\leq n)$.
@@ 4.1 Usual $C$-Gorenstein subcategories
A Wakamatsu tilting module $_{R}C$ with $S=\mathop{\rm End}\nolimits(_{R}C)$ (equivalently, a semidualizing bimodule $_{R}C_{S}$}) is called faithful if the following conditions are satisfied: (i) if $M\in\mathop{\rm Mod}\nolimits R$ satisfying $M_{*}=0$, then $M=0$; and (ii) if $N\in\mathop{\rm Mod}\nolimits S^{op}$ satisfying $N_{*}=0$, then $N=0$ ([16]). In the following result, the implications $(1.1)\Longleftrightarrow(1.2)\Longleftrightarrow(2.1)\Longleftrightarrow(2.2)$ have been obtained in [22, Theorem 5.4(1)].
@@
      $\mathop{\rm G\text{-}pd}\nolimits_{R}\mathcal{I}(R)\leq n$ and $\mathop{\rm id}\nolimits\mathcal{P}(R)\leq n$.
@@ 2. $(b)$
It yields $(4)\Longrightarrow(3)$ immediately. In addition, it also yields $(1.4)\Longleftrightarrow(3)\Longleftrightarrow(2.4)$ by Corollary 4.7. ∎
@@
Now set $R^{\prime}:=R^{op}$. Then $R^{\prime}$ is a von Neumann regular ring, and hence $R^{\prime}$ is a left and right coherent ring and $\mathop{\rm FP\text{-}id}\nolimits_{R^{\prime}}R^{\prime}=\mathop{\rm FP\text{-}id}\nolimits_{{R^{\prime}}^{op}}R^{\prime}=0$. By [27, Theorem 5], we have $\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{FI}(R^{\prime})$. In addition, we have $\mathop{\rm id}\nolimits_{R^{\prime}}R^{\prime}=0$ and $\mathop{\rm id}\nolimits_{{R^{\prime}}^{op}}R^{\prime}\neq 0$ by the above argument, which implies that $R^{\prime}$ is not semisimple, and thus $\mathcal{P}(R^{\prime})\subsetneqq\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{F}(R^{\prime})$ and $\mathcal{I}(R^{\prime})\subsetneqq\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{FI}(R^{\prime})$. In this case, we have $(3)\nRightarrow(1.4)$ and $(3)\nRightarrow(2.4)$ in Proposition 4.11.
@@ 4.2 Auslander and Bass classes
Note that $\mathcal{P}_{C}(S^{op})=\mathop{\rm Add}\nolimits C_{S}$ by [26, Proposition 2.4(a)]. If $\mathcal{P}_{C}(S^{op})\text{-}\mathop{\rm id}\nolimits S=n<\infty$, then it is easy to see that $\mathcal{P}_{C}(S^{op})\text{-}\mathop{\rm id}\nolimits F\leq n$ for any free right $S$-module $F$. It follows from the symmetric version of [37, Lemma 4.6] that $\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits P\leq n$ for any projective right $S$-module $P$. This shows $\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits\mathcal{P}(S^{op})\leq\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits S$, and hence
@@
By [38, Lemma 4.8] and the symmetric version of Proposition 4.1(1) and, we have
@@
    It is natural to ask whether the conditions in Theorem 4.6(i) are left-right symmetric. In general, the answer is no. Otherwise, if the conditions in Theorem 4.6(i) are left-right symmetric, then so are those in Corollary 4.7(i). It yields
@@
Recall that a ring is called semilocal if $R/J(R)$ is a semisimple ring, where $J(R)$ is the Jacobson radical of $R$.

### 11 Mar 2026 v2

@@ 1 Introduction
In Section 4, we apply the results obtained in Section 3 to module categories. The study of generalized tilting modules, which were usually called Wakamatsu tilting modules, was initiated by Wakamatsu [40]. It follows from [42, Corollary 3.2] that $_{R}C$} with $S=\mathop{\rm End}\nolimits(_{R}C)$ is Wakamatsu tilting if and only if $C_{S}$} is Wakamatsu tilting with $R=\mathop{\rm End}\nolimits(C_{S})$, and if and only if $_{R}C_{S}$ is semidualizing bimodule in the sense of [16]. Many authors have studied the properties of Wakamatsu tilting modules and semidualizing bimodules and related modules, see [1, 3], [12], [15, 16], [19][22], [26], [28], [35][42] and references therein.
Let $R$ be an arbitrary ring and $_{R}C$ a Wakamatsu tilting module with $S=\mathop{\rm End}\nolimits(_{R}C)$. We use $\mathcal{P}_{C}(R)$ and $\mathcal{F}_{C}(R)$ to denote the classes consisting of $C$-projective and $C$-flat left $R$-modules, respectively, and use $\mathcal{I}_{C}(S)$ and $\mathcal{FI}_{C}(S)$ to denote the classes consisting of $C$-injective and $C$-FP-injective left $S$-modules, respectively (see Section 2.2 for the definitions of these modules). As an application of Theorem 1.1, we get the following result.
@@
      The category of left $R$-modules is right $n$-$(\mathcal{P}_{C}(R),\mathcal{P}_{C}(R))$-Gorenstein.
@@
      The category of left $S$-modules is left $n$-$(\mathcal{I}_{C}(S),\mathcal{I}_{C}(S))$-Gorenstein.
@@
The Auslander class $\mathcal{A}_{C}(S)$ and Bass class $\mathcal{B}_{C}(R)$ with respect to a semidualizing bimodule $_{R}C_{S}$} (equivalently, a Wakamatsu tilting module $_{R}C$ with $S=\mathop{\rm End}\nolimits(_{R}C)$) originated from [12] are now collectively known as Foxby classes, and they are linked together by Foxby equivalence [12, 16] and duality pairs [20]. These two classes of modules play a crucial role in studying the homological theory related to semidualizing bimodules and Wakamatsu tilting modules, see [1, 12, 15, 16], [19][22], [26, 28], [35][42] and references therein.
@@
    The category of left $R$-modules is left $n$-$(\mathcal{P}_{C}(R),\mathcal{P}_{C}(R))$-Gorenstein.
@@
    The category of left $S$-modules is right $n$-$(\mathcal{I}_{C}(S),\mathcal{I}_{C}(S))$-Gorenstein.
@@ 2.1 Relative homological dimensions and one-sided Gorenstein subcategories
$\text{in}\ \mathscr{A}\ \text{with all}\ X_{i}\ \text{in}\ \mathscr{X}\},$
@@
$\text{in}\ \mathscr{A}\ \text{with all}\ X^{i}\ \text{in}\ \mathscr{X}\},$
@@ 2.2 Wakamatsu tilting modules
Generalized tilting modules are usually called Wakamatsu tilting modules, see [3, 28]. Let $C$ be a left $R$-module and $S=\mathop{\rm End}\nolimits(_{R}C)$. It follows from [42, Corollary 3.2] that $_{R}C$} is Wakamatsu tilting if and only if $C_{S}$} is Wakamatsu tilting with $R=\mathop{\rm End}\nolimits(C_{S})$, and if and only if $_{R}C_{S}$ is semidualizing bimodule in the sense of [1, 16]. Typical examples of Wakamatsu tilting modules include the free module of rank one and the dualizing module over a Cohen-Macaulay local ring. For more examples of Wakamatsu tilting modules, the reader is referred to [16, 37, 41].
@@ 2. $(b)$
    in $\mathscr{A}$ with all $U_{i}$ in $\mathscr{U}$.
@@
in $\mathscr{A}$ with all $U_{i}$ in $\mathscr{U}$ is $\mathop{\rm Hom}\nolimits_{\mathscr{A}}(\mathscr{U},-)$-exact, and the assertion follows.
@@
    in $\mathscr{A}$ with all $E^{i}$ in $\mathscr{E}$.
@@ 3.2 One-sided Gorenstein categories
in $\mathscr{A}$ with all $G_{i}$ in $r\mathcal{G}(\mathscr{C},\mathscr{D})$. For any $D\in\mathscr{D}$, applying the functor $\mathop{\rm Hom}\nolimits_{\mathscr{A}}(-,D)$ to the above exact sequence yields
@@
It is easy to get $(4)\Longrightarrow(1)$ and $(5)\Longrightarrow(3)$.
@@
    $\mathscr{A}$ is right $n$-$\mathscr{C}$-Gorenstein ((that is, $\mathscr{C}$-$\mathop{\rm pd}\nolimits\mathcal{I}(\mathscr{A})\leq n$ and $\mathop{\rm id}\nolimits\mathscr{C}\leq n)$.
@@ 4.1 Usual $C$-Gorenstein subcategories
A Wakamatsu tilting module $_{R}C$ with $S=\mathop{\rm End}\nolimits(_{R}C)$ (equivalently, a semidualizing bimodule $_{R}C_{S}$}) is called faithful if the following conditions are satisfied: (i) if $M\in\mathop{\rm Mod}\nolimits R$ satisfies $M_{*}=0$, then $M=0$; and (ii) if $N\in\mathop{\rm Mod}\nolimits S^{op}$ satisfies $N_{*}=0$, then $N=0$ ([16]). In the following result, the implications $(1.1)\Longleftrightarrow(1.2)\Longleftrightarrow(2.1)\Longleftrightarrow(2.2)$ have been obtained in [22, Theorem 5.4(1)].
@@
      $\mathop{\rm G\text{-}pd}\nolimits\mathcal{I}(R)\leq n$ and $\mathop{\rm id}\nolimits\mathcal{P}(R)\leq n$.
@@ 2. $(b)$
It yields $(4)\Longrightarrow(3)$ immediately. In addition, it yields $(1.4)\Longleftrightarrow(3)\Longleftrightarrow(2.4)$ by Corollary 4.7. ∎
@@
Now set $R^{\prime}:=R^{op}$. Then $R^{\prime}$ is also a von Neumann regular ring, and hence $R^{\prime}$ is a left and right coherent ring and $\mathop{\rm FP\text{-}id}\nolimits_{R^{\prime}}R^{\prime}=\mathop{\rm FP\text{-}id}\nolimits_{{R^{\prime}}^{op}}R^{\prime}=0$. By [27, Theorem 5], we have $\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{FI}(R^{\prime})$. In addition, we have $\mathop{\rm id}\nolimits_{R^{\prime}}R^{\prime}=0$ and $\mathop{\rm id}\nolimits_{{R^{\prime}}^{op}}R^{\prime}\neq 0$ by the above argument, which implies that $R^{\prime}$ is not semisimple, and thus $\mathcal{P}(R^{\prime})\subsetneqq\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{F}(R^{\prime})$ and $\mathcal{I}(R^{\prime})\subsetneqq\mathop{\rm Mod}\nolimits R^{\prime}=\mathcal{FI}(R^{\prime})$. In this case, we have $(3)\nRightarrow(1.4)$ and $(3)\nRightarrow(2.4)$ in Proposition 4.11.
@@ 4.2 Auslander and Bass classes
Note that $\mathcal{P}_{C}(S^{op})=\mathop{\rm Add}\nolimits C_{S}$ by the symmetric version of [26, Proposition 2.4(a)]. If $\mathcal{P}_{C}(S^{op})\text{-}\mathop{\rm id}\nolimits S=n<\infty$, then it is easy to see that $\mathcal{P}_{C}(S^{op})\text{-}\mathop{\rm id}\nolimits F\leq n$ for any free right $S$-module $F$. It follows from the symmetric version of [37, Lemma 4.6] that $\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits P\leq n$ for any projective right $S$-module $P$. This shows $\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits\mathcal{P}(S^{op})\leq\mathcal{P}_{C}(S^{op})$-$\mathop{\rm id}\nolimits S$, and hence
@@
By [38, Lemma 4.8] and the symmetric versions of [20, Proposition 3.2] and Proposition 4.1(1), we have
@@
    It is natural to ask whether the conditions in Theorem 4.6(i) are left-right symmetric. In general, the answer is negative. Otherwise, if the conditions in Theorem 4.6(i) are left-right symmetric, then so are those in Corollary 4.7(i). It yields
@@
Recall that a ring $R$ is called semilocal if $R/J(R)$ is a semisimple ring, where $J(R)$ is the Jacobson radical of $R$.